C=85+0.5Yd

I=85 C=60

net tax T=40+0.25Y

What is equilibrium of National income


1
Expert's answer
2020-09-08T10:14:01-0400

To find the equilibrium of National Income, Aggregate Demand =Aggregate Supply.


AE=C+I

=85+0.5(Y-[40+0.25Y])+60+85


=230+(0.5[Y–40– 0.25Y])


=230+0.5(–0.75Y–40)


=230–0.375Y–20 =0


=210=0.375Y


Y=560